2. Two small beads having positive charges q, 3q and q2 q are fixed at the opposite ends of a horizontal insulating rod of length d 1.50 m. The bead with charge q1 is at the left end and the bead with charge q2 is at the right end of the rod. A third small charged bead is free to slide on the rod. Find the position x where the third bead is in equilibrium 91 42
Answer #1

here,

q1 = 3q , q2 = q

d = 1.5 m

the magnitude of feild on q , E = K * q1 * q /x^2 - K * q2 * q /(d - x)^2

as the net electric feild is zero

0 = K * q1 * q /x^2 - K * q2 * q /(d - x)^2

0 = 1 /x^2 - 3 /(1.5 - x)^2

solving for x

x = 0.55 m

the position x is 0.55 m